# MAC-VO: Metrics-aware Covariance for Learning-based Stereo Visual Odometry mac-vo.github.io

### v1

@@ B-B Robustness Analysis
TABLE X: Robustness of systems on TartanAir v2 Hard test dataset demonstrate by the average variance of relative translation and rotation error.
@@ B-C Additional Ablation Study
TABLE XI: Performance comparison of different ablation setups on the TartanAir v2 Hard Dataset.
@@
TABLE XII: Performance comparison of different ablation setups on the TartanAir v2 Easy Dataset.
@@ B-E Network and Training Details
In the uncertainty-aware matching network, we employ a flow decoder similar to Flowformer to obtain $\Delta f$. In each iteration, the flow $f_{x}$ is updated by $f_{x}\leftarrow f_{x}+\delta f$. With the shared features, we train a new decoder with ConvGRU layers to estimate the uncertainty updates $\Delta\Sigma$. To stabilize the training, we employ $\exp$ as the activation function of the covariance decoder.
For each pixel $x$, we extract the local cost-map patch $q_{x}$ from the cost map, and the context feature $t_{x}$ from the context encoder. We encode the $q_{x}$ using a transformer FFN. Given the 2D position $p=x+f_{x}$, we encode it into positional embedding $\text{PE}(p)$. We aggregate these features using CNN encoder ME to generate local motion features ${I_{x}}$. Utilizing the GMA module, the network’s global motion features $G_{x}$ are acquired from the current motion features $\text{Att}_{f}(t_{x})$ and ${I_{x}}$. Subsequently, the shared motion features $m_{x}$ is obtained by concatenating the $G_{x}$, $I_{x}$ and $\text{Att}(t_{x})$, where $\text{Att}(t_{x})$ is the attention of the context features. To estimate the flow update $\Delta f$ and $\Delta\Sigma$, we us the ConvGRU $\Sigma_{\pi}(m_{x})$
@@ B-F GPU memory & Parameters
As shown in Table. XIII, we use 4.20GB of GPU memory, which is 6.7 times smaller than DROID-SLAM. This is because we utilize the sparse features for back-end optimization, reducing the requirements for the GPU memory.
TABLE XIII: The GPU memory consumption and the number of parameters during testing
@@ Appendix C Runtime Analysis
The runtime analysis use the platform with AMD Ryzen 9 5950X CPU and NVIDIA 3090 Ti GPU. To speed up the frontend network, we utilize CUDAGraph [^1] https://docs.nvidia.com/cuda/cuda-c-programming-guide/index.html#cuda-graphs (CUDAGraph), TensorRT [^2] https://developer.nvidia.com/tensorrt (TRT), and multi-processing (MP) to accelerate the original network (Raw). In our multiprocessing setup, as shown in Fig. B, a new CPU process is initiated to run the pose graph optimization in parallel with the front-end network, thereby maximizing GPU utilization.
We also introduce a fast mode (MAC-VO Fast) which utilizes half-precision (float16) number in the network inference to enhance computational efficiency. This mode also speeds up the memory decoder network by reducing the number of iterative updates from 12 to 4. The fast mode performs 10.5 fps (frames per second) with 70% of the performance of the original MAC-VO.

### v2

@@ B-B Robustness Analysis
See Table IX for data.
TABLE IX: Robustness of systems on TartanAir v2 Hard test dataset demonstrate by the average variance of relative translation and rotation error.
@@ B-C Additional Ablation Study
See Table X and Table XI for data.
TABLE X: Performance comparison of different ablation setups on the TartanAir v2 Hard Dataset.
@@
TABLE XI: Performance comparison of different ablation setups on the TartanAir v2 Easy Dataset.
@@ B-E Network and Training Details
In the uncertainty-aware matching network, we employ a flow decoder similar to Flowformer to obtain $\Delta f$. In each iteration, the flow $f_{x}$ is updated by $f_{x}\leftarrow f_{x}+\delta f$. With the shared features, we finetune a new decoder with ConvGRU layers to estimate the uncertainty updates $\Delta\sigma$. In each iteration, the matching uncertainty is updated by $\sigma\leftarrow\sigma\times\Delta\sigma$.
For each pixel $x$, we extract the local cost-map patch $q_{x}$ from the cost map, and the context feature $t_{x}$ from the context encoder. We encode the $q_{x}$ using a transformer FFN. Given the 2D position $p=x+f_{x}$, we encode it into positional embedding $\text{PE}(p)$. We aggregate these features using CNN encoder ME to generate local motion features ${I_{x}}$. Utilizing the GMA module, the network’s global motion features $G_{x}$ are acquired from the current motion features $\text{Att}_{f}(t_{x})$ and ${I_{x}}$. Subsequently, the shared motion features $m_{x}$ is obtained by concatenating the $G_{x}$, $I_{x}$ and $\text{Att}(t_{x})$, where $\text{Att}(t_{x})$ is the attention of the context features. To estimate the flow update $\Delta f$ and $\Delta\Sigma$, we use the ConvGRU $\Sigma_{\pi}(m_{x})$
@@ B-F GPU memory & Parameters
As shown in Table XII, we use 4.20GB of GPU memory, which is 6.7 times smaller than DROID-SLAM. This is because we utilize the sparse features for back-end optimization, reducing the requirements for the GPU memory.
TABLE XII: The GPU memory consumption and the number of parameters during testing
@@ Appendix C Runtime Analysis
The runtime analysis use the platform with AMD Ryzen 9 5950X CPU and NVIDIA 3090 Ti GPU. To speed up the frontend network, we utilize TensorRT [^1] https://developer.nvidia.com/tensorrt (TRT), and multi-processing (MP) to accelerate the original network (Raw). In our multiprocessing setup, as shown in Fig. 11, a new CPU process is initiated to run the pose graph optimization in parallel with the front-end network, thereby maximizing GPU utilization.
We also introduce a fast mode (MAC-VO Fast) which utilizes half-precision (bfloat16) number in the network inference to enhance computational efficiency. This mode also speeds up the memory decoder network by reducing the number of iterative updates from 12 to 4. The fast mode performs 10.5 fps (frames per second) with 70% of the performance of the original MAC-VO.